DISPERSED FLUORESCENCE SPECTROSCOPY OF CHD_{2}O: A PSEUDO-JAHN-TELLER MOLECULE WITH INTERESTING EMISSION SPECTRA

VADIM STAKHURSKY, \underline{ILIAS SIOUTIS}, VLADIMIR LOZOVSKY^{a}, TERRY A. MILLER, and C. BRADLEY MOORE^{b}, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Methoxy radical (CH_{3}O) is an important intermediate in the combustion of hydrocarbon fuels. It is also an interesting target for theoretical studies, since the doubly degenerate ground X̃^{2}E electronic state exhibits Jahn-Teller (JT) activity. Upon asymmetric deuteration the electronic state splits into two pseudo-degenerate states. To address the issue of the reduction of symmetry in JT active systems we report the laser excited, dispersed fluorescence (DF) spectra of CHD_{2}O. A comprehensive model Hamiltonian which includes linear and quadratic JT coupling terms is developed and used to assign ground state vibronic levels up to 2000 cm^{−1} above the vibrationless level. The experimental observations and intensity simulations for 5 DF traces, from the 3^{4}, 3^{2}5^{′1}, 3^{2}5^{′′1}, 3^{2}6^{′1}, 3^{2}6^{′′1} vibrational levels of the Ã^{2}A′ state, will be presented.

aDeceased

bCurrent address: Northwestern University, 633 Clark Street, Evanston, IL 60208-1108